OpenStudy (anonymous):

When a polynomial P (x) is divided by x +3, the remainder is 2. Which point must be on the graph of the corresponding function y = P (x)?

OpenStudy (anonymous):

A. −3, −2 B. −3, 0 C. −3, 2 D. 3, 2

OpenStudy (anonymous):

I tried reverse engineering the question to get me x+5 for P(x) but that doesn't work for any of the points.

OpenStudy (anonymous):

\[\frac{P(x)}{x+3}=Q(x)+\frac{2}{x+3}\] Because there's a remainder when you divide, you know that \((x+3)\) is not a factor of \(P(x)\), and so \(x=-3\) is not a root. This means you can eliminate (b) \((-3,0)\). From the Remainder Theorem we also have \[P(x)=(x+3)Q(x)+2\] Plugging in \(x=-3\), we have \[P(-3)=(-3+3)Q(-3)+2=2\] which is the point \((-3,2)\).